# Aerial Vision-Language Navigation with a Unified Framework for Spatial, Temporal and Embodied Reasoning

### v1

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) aims to enable unmanned aerial vehicles (UAVs) to interpret natural language instructions and navigate complex urban environments using onboard visual observation. This task holds promise for real-world applications such as low-altitude inspection, search-and-rescue, and autonomous aerial delivery. Existing methods often rely on panoramic images, depth inputs, or odometry to support spatial reasoning and action planning. These requirements increase system cost and integration complexity, thus hindering practical deployment for lightweight UAVs. We present a unified aerial VLN framework that operates solely on egocentric monocular RGB observations and natural language instructions. The model formulates navigation as a next-token prediction problem, jointly optimizing spatial perception, trajectory reasoning, and action prediction through prompt-guided multi-task learning. Moreover, we propose a keyframe selection strategy to reduce visual redundancy by retaining semantically informative frames, along with an action merging and label reweighting mechanism that mitigates long-tailed supervision imbalance and facilitates stable multi-task co-training. Extensive experiments on the Aerial VLN benchmark validate the effectiveness of our method. Under the challenging monocular RGB-only setting, our model achieves strong results across both seen and unseen environments. It significantly outperforms existing RGB-only baselines and narrows the performance gap with state-of-the-art panoramic RGB-D counterparts. Comprehensive ablation studies further demonstrate the contribution of our task design and architectural choices.
@@ I Introduction
Unmanned Aerial Vehicle (UAV) has become an indispensable tool in modern remote sensing applications, playing a central role in infrastructure inspection, environmental monitoring, and emergency response [khan2022uavsurvey, yao2024autonomous]. Previous research has largely focused on passive perception tasks, including object detection [xi2024dronedetection1, wu2025dronedetection2] and tracking [yuan2025dronevideo, xiao2025uav] from aerial images or videos, without interaction with the world. In contrast, aerial navigation tasks require the drone to perceive, reason, and act in dynamic environments. Recently, aerial Vision-and-Language Navigation (VLN) [liu2023aerialvln] has emerged as a new paradigm, where drones follow high-level language instructions to navigate the destination through 3D outdoor environments. By leveraging natural language as a human-centric interface, aerial VLN significantly reduces the reliance on expert pilots, lowers the barrier of human-UAV interaction, and enables intuitive task specification in high-stakes scenarios.
Recent works have explored the design of benchmarks and datasets to facilitate research in aerial vision-and-language navigation. AVDN [fan2023AVDN] first proposed a dialogue-based setting involving asynchronous interactions between a human commander and a UAV agent. AerialVLN [liu2023aerialvln] introduced high-fidelity city-scale simulations with diverse human-annotated trajectories. CityNav[lee2025citynavdataset] extended this line by leveraging real-world urban reconstructions. OpenUAV [wangopenuav] formulated aerial VLN as full-trajectory prediction with human-in-the-loop evaluation, and OpenFly[gao2025openfly] scaled up scene and instruction diversity via automatic generation. AeroDuo [wu2025aeroduo] explores the collaborative setting with multi-agent instruction following with altitude-aware role assignment. Moreover, CityEQA [zhao2025cityeqa] and 3D Open-EQA [zhang2025open3dvqa] have proposed embodied question answering benchmarks to assess model’s perception and reasoning capabilities from aerial perspective. However, as illustrated in Fig. 1, aerial VLN requires an agent to continuously integrate current observation, trajectory history, and high-level instruction. Concretely, the agent must recognize location from egocentric views, assess task progress relative to instruction, and infer the next action consistent with the intended route. Aerial VLN presents unique challenges across three core aspects.
@@
Massive endeavors have attempted to alleviate these challenges. A line of work methods leverages pre-trained vision-language models (VLMs) to enable zero-shot instruction following. STMR [gao2024aerialstmr] encodes current egocentric view, drone pose, and subgoal cues into a matrix-form prompt, allowing the language model to reason over spatial structures and landmarks. CityNavAgent [zhang2025citynavagent] enhances semantic grounding by constructing symbolic 2D landmark maps from real aerial imagery combined with object detection. In contrast, other line of work aims to directly align visual observations with language instructions through joint training. LAG [liu2023aerialvln] combines a look-ahead guidance mechanism to Cross-Modal Attention method [krantz2020cma]. CityNav [lee2025citynavdataset] introduces geographic semantic map (GSM), which integrates map data retrieved from OpenStreetMap with the pose of drones. Grid-based View Selection [zhao2025aerialviewselection] reformulates the navigation task as a coupled horizontal–vertical view selection problem, integrating bird’s-eye fusion and cross-modal alignment. While these approaches have achieved promising performance, many of them rely on auxiliary inputs such as panoramic cameras, depth sensors, odometry or maps. These modalities lead to significant hardware cost and complexity. These design choices hinder generalizability and make them less suitable for real-world UAV deployment.
In this paper, we aim to develop a unified framework for aerial VLN, which relies solely on onboard egocentric RGB observations and natural language instructions. Without the need of any panoramic images, depth information, odometry and maps, our method directly predicts the next action step-by-step, bridging language semantics with visual context. Under the Next-Token Prediction (NTP) paradigm, we jointly models spatial perception, trajectory reasoning, and high-level action prediction within a single vision-language backbone.Task-specific prompts are introduced to steer learning towards distinct reasoning objectives while preserving architectural elegance and training efficiency. To mitigate redundancy in long-horizon video trajectories, we adopt an egocentric keyframe selection mechanism that retains informative frames and significantly reduces the computational burden caused by massive visual tokens. Furthermore, we design an action-merge and label reweighting strategy to alleviate label imbalance and facilitate multi-task co-training. All tasks are co-optimized under a unified vision-language interface, encouraging the model to satisfied the nature of aerial VLN, which demands precise spatial grounding, temporal abstraction, and embodied control. Experimental results demonstrate that our method con sistently outperforms baseline models. Our contributions can be summarized as follows:
@@
    We develop an egocentric aerial VLN framework that relies solely on monocular front-view RGB observations and natural language instructions, removing the need for panoramic inputs, depth sensors, odometry, or pre-built maps. The proposed model directly maps textual instructions and egocentric visual observations into executable navigation actions.
@@
    We formulate aerial VLN as a unified next-token prediction problem, where spatial perception, trajectory reasoning, and action generation are jointly optimized under a unified vision-language interface via task-specific prompts. Moreover, we propsed a keyframe selection strategy to reduce visual redundancy, along with an action-merge and label reweighting mechanism to mitigate label imbalance.
@@
    We conduct extensive experiments on standard benchmarks, achieving state-of-the-art performance and validating the effectiveness of each proposed component through comprehensive ablation studies.
@@ II Related Works
II-A Aerial Navigation
Aerial navigation has attracted growing attention, with early efforts focusing on conventional tasks such as obstacle avoidance [lindqvist2021obstacleavoidance, zhang2025obstacleavoidancenature] and object tracking [wang2019targettracking, savkin2023effective] based on onboard cameras or GPS signals. Recent advances explore language-guided aerial navigation, aiming to empower UAVs to interpret and execute natural language instructions in large-scale, open-world scenarios [liu2023aerialvln]. This task introduces additional challenges beyond visual navigation, such as grounding language instructions into visual perception and continuous control, performing temporal reasoning, and aligning multimodal information for robust decision-making. AVDN [fan2023AVDN] introduces a dialogue-based UAV-VLN setting with asynchronous commander–follower interactions. AerialVLN [liu2023aerialvln] collects diverse human-piloted trajectories paired with spatially grounded instructions in synthetic outdoor cities, while CityNav [lee2025citynavdataset] leverages real-world 3D urban reconstruction. OpenUAV [wangopenuav] formulates UAV-VLN as full-trajectory prediction and incorporates human-in-the-loop evaluation. OpenFly [gao2025openfly] scales up dataset diversity through automatic generation over varied high-fidelity virtual environments.
Early approaches to aerial VLN adapted models from ground-based navigation but often struggled with dynamic viewpoints and large-scale outdoor complexity. More recently, with the rise of large language models (LLMs) and vision language models (VLMs), several works have explored training-free paradigms that exploit their visual reasoning and visual grounding abilities for zero-shot aerial VLN [liu2024framework, zhang2024demo, yumeng2025empowering]. The line for this work converts visual observations into text prompts, in the form of topological graphs, symbolic maps, or structured semantic matrices and rely on LLMs to infer the next action or waypoint. Among them, STMR [gao2024aerialstmr] introduces a Semantic-Topo-Metric Representation that encodes current view, pose, and subgoal cues into matrix-form prompts for spatial reasoning over landmarks and directions. CityNavAgent [zhang2025citynavagent] further enhances grounding via symbolic 2D landmark maps derived from real aerial imagery and spatial detection. In contrast, learning-based models directly align visual observations with language through end-to-end training. AeiralVLN [liu2023aerialvln]proposed the Look-ahead Guidance (LAG) strategy to improve CMA. Grid-based VS [zhao2025aerialviewselection] reformulates aerial VLN as a grid-based view selection problem with coupled vertical-horizontal action modeling, bird’s-eye history fusion, and cross-modal alignment. More related to ours method, OpenFly [gao2025openfly] introduces keyframe-aware VLN and adaptive token merging to better handle long-horizion trajectories.
II-B Vision-and-Language Navigation
@@
Vision-and-language navigation (VLN) tasks require an agent to follow natural language instructions using visual observations to achieve a target location [wu2024vln_survey]. Early VLN benchmarks formulated navigation on a predefined graph of panoramic viewpoints, where the agent selects discrete actions to move between connected nodes [anderson2018r2r, ku2020rxr]. To move beyond the discrete setting, VLN-CE [krantz2020vlnce] replaced the navigation graph with free movement in simulation, where the agent executes mid-level actions, such as moving forward and turning. A variety of learning-based methods have been proposed under this setting, including cross-modal attention [hong2021vln_bert], memory-augmented agents [wang2023gridmm], and planning-based approaches [an2024etpnav, wang2023dreamwalker]. A growing line of work has focused on training vision-language models that directly map RGB observations and instructions to navigation actions. NaVid [zhang2024navid] demonstrates that a vision-language model trained on large-scale continuous trajectories can effectively use monocular RGB video to model temporal context and predict actions, achieving strong performance. NaVILA [cheng2025navila] introduces a two-stage framework that maps visual observations and instructions into mid-level language actions, which are then executed by a learned locomotion policy, enabling instruction-conditioned control in low-level continuous environments. MonoDream [wang2025monodream] introduces auxiliary representation learning tasks that train a monocular agent to align visual semantics and action intent by predicting latent panoramic features. Uni-NaVid [zhang2024uninavid] unifies diverse navigation tasks within a single end-to-end framework by learning from multi-task RGB video data. VLN-R1 [qi2025vln] combines supervised and reinforcement fine-tuning to enhance temporal reasoning and multi-step action planning.
@@
While these approaches have made substantial progress, they are primarily developed for ground-based agents and remain limited in capturing the challenges inherent to aerial-based navigation, such as viewpoint variability, long-range flight paths, and complex 3D outdoor scenes.
@@ III Method
III-A Problem Formulation
[…]
Figure 2: Trajectory statistics before and after data preprocessing, showing that action merging and keyframe selection yield a richer action space and more compact navigation sequences. For brevity, we use MF for move forward etc.
@@ III-B Data Preprocessing
Keyframe Selection. To further suppress temporal visual redundancy, we extract keyframes at the boundaries of the merged action segments. This simple yet effective strategy is motivated by the observation that abrupt changes in the agent’s motion—such as transitioning from straight movement to a turning action—are often caused by the appearance of landmarks. These turning points naturally delimit semantically meaningful transitions in the underlying trajectory. By selecting the frames at these boundaries as keyframes, we obtain a compact set of observations that preserves landmarks and other salient visual cues while filtering out redundant intermediate views. This yields a more informative and temporally structured visual stream for downstream navigation learning.
Figure 3: Overview of our framework. Given egocentric keyframes selected from the onboard video stream, our model first encodes the visual observations through a vision encoder and a MLP projector to obtain visual tokens, while language instructions are processed by a text tokenizer. The unified multimodal tokens are then fed into a large language model that is jointly trained on three complementary tasks: (i) Spatial Perception, which queries the current scene; (ii) Trajectory Reasoning, which summarizes historical motion and infers the agent’s navigational context; and (iii) Embodied Navigation, which predicts high-level action commands. The predicted textual action is further parsed and decomposed into a sequence of predefined motion primitives for execution in the physical environment.

### v2

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) aims to enable unmanned aerial vehicles (UAVs) to interpret natural language instructions and navigate complex urban environments using onboard visual observation. This task holds promise for real-world applications such as low-altitude inspection, search-and-rescue, and autonomous aerial delivery. Existing methods often rely on panoramic images, depth information, or odometry to support spatial reasoning and action planning. These requirements increase system cost and integration complexity, thereby hindering the practical deployment of lightweight UAVs. In this work, we present a unified aerial VLN framework that operates solely on egocentric monocular RGB observations and natural language instructions. The model formulates navigation as a next-token prediction problem, jointly optimizing spatial perception, trajectory reasoning, and action prediction through prompt-guided multi-task learning. Moreover, we propose a keyframe selection strategy to reduce visual redundancy by retaining semantically informative frames, along with an action merging and label reweighting mechanism that mitigates long-tailed supervision imbalance and facilitates stable multi-task co-training. Extensive experiments on the AerialVLN and OpenFly benchmark validate the effectiveness of our method. Under the challenging monocular RGB-only setting, our model achieves strong results across both seen and unseen environments. It significantly outperforms existing monocular baselines and narrows the performance gap with state-of-the-art panoramic RGB-D counterparts. Comprehensive ablation studies further validate the contribution of our architectural design and choices.
@@ I Introduction
Unmanned Aerial Vehicle (UAV) has become an indispensable tool in modern remote sensing applications, playing a central role in infrastructure inspection, environmental monitoring, and emergency response [17]. Previous research has largely focused on passive perception tasks, including object detection [21, 9] and tracking [7, 35] from aerial images or videos, without interaction with the world. In contrast, aerial navigation tasks require the drone to perceive, reason, and act in dynamic environments. Recently, aerial Vision-and-Language Navigation (VLN) [25] has emerged as a new paradigm, where drones follow high-level language instructions to navigate the destination through 3D outdoor environments. By leveraging natural language as a human-centric interface, aerial VLN significantly reduces the reliance on expert pilots, lowers the barrier of human-UAV interaction, and enables intuitive task specification in high-stakes scenarios.
Recent works have explored the design of benchmarks and datasets to facilitate research in aerial vision-and-language navigation. AVDN [11] first proposed a dialogue-based setting involving asynchronous interactions between a human commander and a UAV agent. AerialVLN [25] introduced high-fidelity city-scale simulations with diverse human-annotated trajectories. CityNav[20] extended this line by leveraging real-world urban reconstructions. OpenUAV [33] formulated aerial VLN as full-trajectory prediction with human-in-the-loop evaluation, and OpenFly [12] scaled up scene and instruction diversity via automatic generation. AeroDuo [36] explores the collaborative setting with multi-agent instruction following with altitude-aware role assignment. Moreover, CityEQA [49] and 3D Open-EQA [46] have proposed embodied question answering benchmarks to assess model’s perception and reasoning capabilities from an aerial perspective. However, as illustrated in Fig. 1, aerial VLN requires an agent to continuously integrate current observation, trajectory history, and high-level instruction. Concretely, the agent must recognize location from egocentric views, assess task progress relative to instruction, and infer the next action consistent with the intended route. Aerial VLN presents unique challenges across three core aspects.
@@
Massive endeavors have attempted to alleviate these challenges. One line of research leverages pre-trained vision language models (VLMs) to enable zero-shot instruction following [13, 45, 39]. These approaches introduce explicit intermediate representations and memory mechanisms to mitigate large-scale spatial reasoning difficulties and preserve long-term spatial context. A large language models are subsequently employed for high-level planning. However, such modular designs decouple perception, representation, and decision-making into separate stages, which may introduce error accumulation and restrict joint optimization across complementary navigation capabilities. Another line of research adopts learning-based end-to-end modeling to directly map visual observations and language instructions to embodied action sequences [25, 48, 20]. These methods enhance cross-modal alignment by incorporating geographic spatial representations or history-aware fusion modules to better utilize spatial cues and history observations. More recent efforts further leverage large vision-language models with strong multimodal capabilities to improve reasoning over aerial trajectories [12, 5, 38]. While these approaches have achieved promising performance, many rely on auxiliary inputs such as panoramic cameras, depth sensors, or odometry, which increase hardware complexity and deployment cost. Moreover, the limited memory context of current models remains a bottleneck, making long-horizon aerial navigation particularly challenging.
To address the aforementioned limitations without introducing additional sensing modalities, while leveraging the strong multimodal capability of vision-language models, we develop a unified framework for aerial VLN that relies solely on onboard egocentric RGB observations and natural language instructions. The proposed framework directly predicts the next navigation action, bridging language semantics with visual context. we formulate navigation under a Next-Token Prediction (NTP) paradigm, jointly modeling spatial perception, trajectory reasoning, and action generation within a single vision-language backbone. This unified formulation enables cross-modal alignment and trajectory-level reasoning to be optimized under a shared objective without architectural modification. Within this framework, Task-specific prompts are introduced to steer learning towards distinct reasoning objectives while preserving architectural simplicity. Furthermore, considering the structural characteristics of aerial navigation, including substantial visual redundancy during continuous flight and imbalance in action distributions, we incorporate an aerial-specific design consisting of keyframe selection, action merging, and label reweighting. These components are integrated into the overall learning framework to improve supervision balance and sequence modeling stability. Experimental results across multiple benchmarks demonstrate that our method consistently outperforms baseline models.
In summary, our paper makes the following contributions:
@@
    Unified Next-Token Formulation: We formulate aerial VLN as a next-token prediction problem, jointly modeling spatial perception, trajectory reasoning, and action generation within a single autoregressive backbone, thereby facilitating tighter cross-modal alignment.
@@
    Prompt-Driven Multi-Task Supervision: We introduce task-specific prompts to incorporate auxiliary objectives for spatial grounding and trajectory reasoning, strengthening the model’s reasoning ability over spatial structure and temporal evolution.
@@
    Aerial-Specific Design: We propose an aerial-specific design and training strategy, including keyframe selection, action merging, and label reweighting, to address visual redundancy and action imbalance inherent in aerial navigation for stable training.
@@ II Related Works
II-A Vision-and-Language Navigation
@@
Vision-and-language navigation (VLN) tasks require an agent to follow natural language instructions using visual observations to achieve a target location [37, 14]. Early VLN benchmarks formulated navigation on a predefined graph of panoramic viewpoints, where the agent selects discrete actions to move between connected nodes [3, 19, 42]. To move beyond the discrete setting, VLN-CE [18] replaced the navigation graph with free movement in simulation, where the agent executes mid-level actions, such as moving forward and turning. A variety of learning-based methods have been proposed under this setting, including cross-modal attention [15], memory-augmented agents [34], and planning-based approaches [2, 30, 6]. A growing line of work has focused on training vision-language models that directly map RGB observations and instructions to navigation actions. NaVid [44] demonstrates that a vision-language model trained on large-scale continuous trajectories can effectively use monocular RGB video to model temporal context and predict actions, achieving strong performance. NaVILA [10] introduces a two-stage framework that maps visual observations and instructions into mid-level language actions, which are then executed by a learned locomotion policy, enabling instruction-conditioned control in low-level continuous environments. MonoDream [32] introduces auxiliary representation learning tasks that train a monocular agent to align visual semantics and action intent by predicting latent panoramic features. Uni-NaVid [43] unifies diverse navigation tasks within a single end-to-end framework by learning from multi-task RGB video data. VLN-R1 [28] combines supervised and reinforcement fine-tuning to enhance temporal reasoning and multi-step action planning.
@@
While these approaches have made substantial progress, they are primarily developed for ground-based agents in indoor environments and typically operate within limited spatial scales. Our work instead focuses on aerial VLN in large-scale urban environment, with inherent challenges, such as viewpoint variability, long-range flight paths, and complex 3D outdoor scenes. we adopt a unified framework and enhance it with prompt-guided auxiliary supervision to strengthen spatial perception and trajectory reasoning over long horizons
II-B Aerial Vision-and-Language Navigation
Aerial navigation has attracted growing attention, with early efforts focusing on conventional tasks such as obstacle avoidance [23, 47] and object tracking [31, 29] based on onboard cameras or GPS signals. Recent advances explore language-guided aerial navigation, aiming to empower UAVs to interpret and execute natural language instructions in large-scale, open-world scenarios [25]. This task introduces additional challenges beyond visual navigation, such as grounding language instructions into visual perception and continuous control, performing temporal reasoning, and aligning multimodal information for robust decision-making.
Research on aerial VLN has gradually evolved with the introduction of dedicated benchmarks and diverse solution paradigms. AVDN [11] introduces a dialogue-based UAV-VLN setting while AerialVLN [25] and CityNav [20] construct large-scale outdoor datasets using synthetic cities and real-world 3D reconstruction, respectively. OpenUAV [33] formulates UAV-VLN as full-trajectory prediction and incorporates human-in-the-loop evaluation. OpenFly [12] scales up dataset diversity through automatic generation over varied high-fidelity virtual environments. Early approaches to aerial VLN adapted models from ground-based navigation, but often struggled with dynamic viewpoints and large-scale outdoor complexity. More recently, with the rise of large language models (LLMs) and vision language models (VLMs), a line of work explores training-free paradigms for zero-shot aerial VLN. These methods convert visual observations into text prompts in the form of topological graphs, symbolic maps, or structured semantic matrices and rely on LLMs to infer the next navigation action or waypoint. Among them, STMR [13] introduces a Semantic-Topo-Metric Representation that encodes current view, pose, and subgoal cues into matrix-form prompts for spatial reasoning over landmarks and directions. CityNavAgent [45] further enhances semantic grounding via symbolic 2D landmark maps derived from real aerial imagery and spatial detection. In contrast, learning-based approaches directly align visual observations and language instructions through end-to-end optimization. LAG [25] introduces a look-ahead guidance mechanism to Cross-Modal Attention method [18]. Grid-based View Selection [48] reformulates aerial VLN as a grid-based view selection problem with coupled vertical-horizontal action modeling, bird’s-eye history fusion, and cross-modal alignment. More closely related to our work, OpenFly [12] proposes keyframe-aware modeling and adaptive token merging to improve efficiency and stability in long-horizon navigation.
However, many existing aerial VLN approaches rely on additional sensory inputs or explicit structural representations to facilitate navigation. Differently, we focus on an egocentric monocular RGB setting and formulate aerial VLN under a unified next-token prediction paradigm. Within this formulation, we incorporate two auxiliary objectives to enhance the scene understanding and temporal reasoning. We further introduce aerial-specific training strategies, including keyframe selection, action merging, and label reweighting, to reduce visual redundancy and mitigate label imbalance.
Figure 2: Trajectory statistics before and after data preprocessing, showing that action merging and keyframe selection yield a richer action space and more compact navigation sequences. For brevity, we use MF for move forward, etc.
@@ III Method
III-A Task Definition
Aerial vision-and-language navigation. The goal of aerial VLN is to learn a policy that enables drones to follow high-level language instructions to navigate through 3D outdoor environments and reach the target location. Specially, an aerial agent is tasked with navigating a continuous 3D space according to a natural language instruction $I=\{w_{1},w_{2},\ldots,w_{N_{I}}\}$, which comprises $N_{I}$ words describing the desired flight route. At each time step $t$, the agent receives a visual trajectory $\tau_{t}=\{x_{1},\ldots,x_{t}\}$, where $x_{t}\in\mathbb{R}^{H\times W\times 3}$ is the first-person RGB image captured by its onboard camera. Conditioned on visual trajectory $\tau_{t}$ and the instruction $I$, the agent selects an action $a_{t}$ from a discrete action set $\mathcal{A}$. Executing $a_{t}$ moves the agent to the next state in the environment, yielding the subsequent image observation $x_{t+1}$. An episode terminates when the agent chooses the stop action or reaches the maximum step limit. A navigation episode is considered successful if the agent’s final position lies within $20$ meters of thedestination. The action space and corresponding motion magnitudes are defined according to the specific environment settings. In AerialVLN environment, $\mathcal{A}=\{\textit{move forward},\allowbreak\textit{turn left},\allowbreak\textit{turn right},\allowbreak\textit{ascend},\allowbreak\textit{descend},\allowbreak\textit{move left},\allowbreak\textit{move right},\allowbreak\textit{stop}\}$. horizontal motions displace the agent by $5$ units, vertical motions by $2$ units, and turning applies a fixed $15^{\circ}$ change in orientation. In OpenFly environment, $\mathcal{A}=\{\textit{move forward},\allowbreak\textit{turn left},\allowbreak\textit{turn right},\allowbreak\textit{ascend},\allowbreak\textit{descend},\allowbreak\textit{ stop}\}$. Horizontal motions displace the agent by $3$ units, vertical motions by $3$ units, and turning applies $30^{\circ}$.
Figure 3: Overview of our framework. Given egocentric keyframes selected from the onboard video stream, our model first encodes the visual observations through a vision encoder and an MLP projector to obtain visual tokens, while language instructions are processed by a text tokenizer. The unified multimodal tokens are then fed into a large language model that is jointly trained on three complementary tasks: (i) Spatial Perception, which queries the current scene; (ii) Trajectory Reasoning, which summarizes historical motion and infers the agent’s navigational context; and (iii) Embodied Navigation, which predicts high-level action commands. The predicted textual action is further parsed and decomposed into a sequence of predefined motion primitives for execution in the physical environment.
@@ III-B Data Preprocessing
Keyframe Selection. To further suppress temporal visual redundancy, we extract keyframes at the boundaries of the merged action segments. This simple yet effective strategy is motivated by the observation that abrupt changes in the agent’s motion—such as transitioning from straight movement to a turning action—are often caused by the appearance of landmarks. These turning points naturally delimit semantically meaningful transitions in the underlying trajectory. By selecting the frames at these boundaries as keyframes, we obtain a compact set of observations that preserves landmarks and other salient visual cues while filtering out redundant intermediate views. This yields a more informative and temporally structured visual stream for downstream navigation learning. Unlike simple uniform sampling [10] or detector-dependent landmark selection [12], our strategy is control-aligned and does not rely on auxiliary perception modules, providing an efficient keyframe selection mechanism.